OpenStudy (anonymous):

Take 2y 2 - 3y - 5 from y 3 - 6y 2 + 5y. Select the correct answer. y3 - 2y2 + 3y + 5 y3 - 4y2 + 2y - 5 y3 - 4y2 + 2y + 5 y3 - 8y2 + 8y + 5

OpenStudy (anonymous):

\[\left[ y3-6y2+5y \right] - \left[ 2y2 - 3y-5 \right] = y3 - 8y2 + 8y + 5\]